OpenStudy (anonymous):

What is the solution to the equation 3 ^1 – 2x = 2 ? x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (anonymous):

confirm your problem. what is show is actually just .5

OpenStudy (anonymous):

What is the solution to the equation 3 ^1 – 2x = 2 ? x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (anonymous):

3 ^1 – 2x = 2 that looks like 3 - 2x = 2 -2x = 2-3 -2x = -1 x =1/2 x =.5

OpenStudy (anonymous):

oh :)